OpenStudy (anonymous):

if = sqrtofx = sqrt of y =4,find y doble prime a. sqrt of y over sqrt of x b. sqrt of x over sqrt of y c. negative sqrt of x over negative sqrt of y d. negative sqrt of y over negative sqrt of x

OpenStudy (anonymous):



OpenStudy (anonymous):

\[find y'' if \sqrt x + \sqrt y =4 \]

OpenStudy (ash2326):

We have \[\sqrt x+\sqrt y=4\] Let's differentiate this with respect to x \[\frac{d}{dx}(\sqrt x+\sqrt y=4)\] we get \[ \frac{1}{2\sqrt x}+ \frac{1}{2\sqrt y} *y'=0\] we have now \[\frac{1}{2\sqrt y} *y'=-\frac{1}{2\sqrt x}\] we have now \[y'=-\frac{\sqrt y}{\sqrt x}\] now let's differentiate this again with respect to x \[y''=-(\frac{\sqrt x (\frac{d}{dx}\sqrt y)-\sqrt y (\frac{d}{dx} \sqrt x) }{x})\] we get now \[y''=-(\frac{\sqrt x (\frac{1}{2\sqrt y} *y')-\sqrt y (\frac{1}{2\sqrt x} )}{x})\] we have \[y'=-\frac{\sqrt y}{\sqrt x}\] so \[y''=-(\frac{\sqrt x (\frac{1}{2\sqrt y} *-\frac{\sqrt y}{\sqrt x})-\sqrt y (\frac{1}{2\sqrt x} )}{x})\] so we get finally \[y''=-(\frac{-\frac{1}{2}-\sqrt y (\frac{1}{2\sqrt x} )}{x})\] we get \[y''=\frac{1+ \frac{\sqrt y}{\sqrt x} }{2x}\]

OpenStudy (anonymous):

in y''' cancel 1 and 2x then the answer is